1.) In a vertical dive, a peregrine falcon can accelerate at 0.60.6 times the free-fall acceleration ?g (that is, at 0.6?0.6g) in reaching a speed of about 104 m/s.104 m/s. If a falcon pulls out of a dive into a circular arc at this speed and can sustain a radial acceleration of 0.6?,0.6g,what is the radius ?R of the turn?

2. )Anne is working on a research project that involves the use of a centrifuge. Her samples must first experience an acceleration of 100?, but then, the acceleration must increase by a factor of six. By how much will the rotational speed have to increase? Express your answer as a fraction of the initial rotation rate.
Answer #1

Solution)

We know, radial acceleration, a= v^2/r

And, a=0.6g= 0.6*9.8=5.88 m/s^2

So, r=v^2/5.88

R= 104^2/5.88=1839.45 m(Ans)

======